Solve the given differential equations:

х 1. y = y(0) = -2 y+x2y 2. 3x²y dx – (x3 + y3)dy = 0 , y(1) = -2
Answer #1

1) Solution y = y+xży; y(0) = -2 Standard form of a first order separable ODE : N(y) y = M(x) х y+x2y Multiply both sides b

= } .l du = { In (u) Substitute back u = 1 + x2 = { In (1 + x²)+c1 = / ydy = + 02 1 ſ ydy == 1+1 == + C2 = { in (1 + x²) + C1